A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1860–1869

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

11th [78, 116, 267, 321]. 12 [116]. 13th [208]. 14th [119]. 1506 [155]. 15th [89, 150, 190, 222, 251, 278, 307, 47, 177]. 16th [278, 320]. 1858 [50]. 1860 [19]. 1861.-Preliminary [62]. 1862 [116]. 1863 [138]. 1864 [177, 182, 168]. 1865 [190, 208]. 1866 [222, 239]. 1867 [267, 273]. 1868 [278, 296, 290]. 1869 [307, 321]. 1980 [23]. 19th [296]. 1st [106].

20th [82, 177]. 22nd [239]. 26th [296, 138]. 27th [23]. 28th [19]. 29th [239, 273]. 2nd [138].

31st [89, 150, 190, 222, 251, 278, 307, 47, 119].

4th [23, 78, 267].

5th [273].

6th [208, 321].

8th [106].

A.D. [155]. Aberdeen [160]. Aboriginal [35]. Abstract [75]. Abstracts [29]. Academical [146]. Accommodation [322]. Account [134, 69, 74]. Accounts [228]. Accumulations [70]. Acquired [203]. Act [16]. Action [224]. Address [15, 202, 290, 105, 182, 168, 262, 320, 312, 232, 19]. Administration [170]. Advancement [23, 78, 106, 138, 177, 202, 208, 239, 267, 296, 290, 105, 168, 262, 312, 232]. Affected [113]. Africa [274]. After [235]. Age [84]. Agents [286]. Agricultural [282, 279, 304, 60, 102, 163, 123]. Agriculture [235]. Alleged [186]. also [203]. Altered [72]. Ameliorations [3]. Amendment [16]. America [281, 145, 113]. Amongst [114, 176]. Analysis [238, 254, 61, 275]. Ancients [53]. Anniversary [8, 89, 150, 190, 222, 251, 278, 307, 47, 119]. Application [18]. Appointments [77]. Approaching [54]. Approximate [212, 200]. April [62, 56]. Arbitration [265]. Argument [51]. Arising [41]. Armies [55, 143]. Arms [204]. Army [133, 67]. Arranged [212, 200]. Ascertained [50]. Ascertainment [18]. Aspect [294]. Association [23, 78, 106, 138, 177, 202, 208, 239, 267, 296, 290, 105, 182, 168, 262, 312, 232, 19]. Associations [71]. Attributed [52]. August [138, 239, 296, 290, 312, 232]. Australia [49, 171]. Australian [49]. Austria [34, 254]. Autumnal [224]. Available [266].

Bank [224]. Banking [246]. Banks [73]. Bath [177, 168]. Baumhauer [136]. be [52, 203]. Bearings [95]. Been [94]. before [16]. being [76, 33]. Bellingham [134]. Benefit [205]. Bengal [91, 245]. Berlin [135, 147, 10]. Between [42, 247]. Birmingham [202, 208, 217, 206, 205, 218]. Births [185, 225]. Body [84, 275]. Books [76]. Borough [217, 84]. Brazil [154]. Britain [40]. British [23, 55, 75, 78, 106, 138, 177, 202, 208, 239, 267, 296, 290, 105, 67, 168, 302, 262, 145, 312, 232, 19, 123, 4]. Budgets [228, 53]. Building [205].

Cambridge [106, 105, 128, 112, 272]. Canada [291]. Carried [31]. Cases [308].

Casualities [152, 200]. Casualties [212, 238, 294]. Catholics [158]. Cattle [292]. Causes [52]. Census [11, 62, 56, 186, 9, 322, 185, 68]. Central [145, 254]. Centuries [76]. Century [114]. Certain [151]. Chair [116]. Charitable [264]. Chie [9]. Chief [170]. Chiefly [52]. Children [40, 214]. China [83]. Christians [192]. Cider [173]. Cider-Truck [173]. City [160]. Civil [77, 258, 113, 187, 293, 303]. Classification [282, 313]. Clearing [196]. Climate [133]. Close [49]. Cloth [69]. Co [71]. Co-operative [71]. Coinage [295]. College [112]. Colonel [116]. Colonial [132]. Colonies [75, 49, 103, 183]. Colony [301]. Coloured [274]. Combinations [244]. Commencement [235]. Commerce [31]. Commercial [52, 183, 145, 33]. Commission [50, 254, 94]. Committee [238]. Communication [116]. Company [291]. Comparative [280, 68, 169, 55]. Compared [133, 206, 170, 286, 323, 123]. Comparison [143]. Competitive [77, 115]. Condition [34, 295, 197, 320, 74, 72, 83, 233]. Conditions [133, 244]. Congress [15, 24, 20, 135, 147, 273, 321, 39, 33]. Connected [137, 313]. Conscription [252]. Consort [15]. Continental [235]. Continuation [32]. Continuous [144]. Contributed [136]. Corn [54]. Cost [303, 143, 216, 162]. Cotton [323, 113, 69]. Council [11, 82, 89, 116, 150, 190, 222, 251, 278, 307, 47, 119]. Counties [265]. Countries [52, 42, 95, 121]. Country [196]. Course [5, 44]. Courts [258]. Crime [206, 286, 308, 257, 104, 161, 171]. Criminal [29, 308]. Cultivation [302]. Currency [193, 295].

Date [76]. Deaths [253, 225]. Debts [95]. Decade [111]. December [116, 82, 89, 150, 190, 222, 251, 278, 307, 47, 119]. Decimal [53]. Decline [314]. Decrease [163].

Defects [151]. Deficiencies [185]. Definitions [191]. Delegates [135, 147]. Delivered [320]. Department [182, 10]. Dependencies [183]. Destitute [40]. Detail [143]. Developing [266]. Development [59]. Dietaries [126]. Direct [41]. Discipline [91, 245]. Discordancy [52]. Discovery [49]. Distressed [97]. Distribution [3]. Districts [2, 145]. down [9]. Dr. [136]. Dundee [267, 262]. Durham [146]. During [114, 111, 323, 66, 44, 121, 5, 76]. Dutch [136].

Earnings [60, 102]. Economic [202, 290, 105, 168, 262, 315, 197, 66, 312, 232, 19, 233]. Economical [95]. Economy [182]. Edibles [155]. Education [50, 128, 112, 86, 266, 248, 50]. Educational [40, 184, 272, 264]. Effect [42]. Effects [77, 67, 41, 70]. Efficiency [20]. Eighteen [84]. Eighth [89, 296, 290]. Elementary [248]. Embroidered [72]. Emigration [103]. Empire [59, 275]. Employed [76, 41]. Employment [291]. Ended [89, 150, 222, 278, 307, 47, 119, 190, 251]. Endowed [112]. Endowments [146]. England [291, 50, 56, 135, 147, 235, 282, 29, 228, 170, 257, 224, 175, 158, 86, 322, 60, 97, 85, 163, 76, 266, 173, 68, 83, 313]. English [258, 16, 185, 143, 216]. Enjoyments [48]. Enterprise [12]. Especial [41]. Especially [21]. Esq [116]. Esq. [19]. Eurasians [159]. Europe [214, 215]. European [133, 280]. Europeans [176]. Examinations [115]. Exchanges [42]. Exercised [94]. Exeter [312]. Exhaustive [18]. Existing [323]. Expenditure [217, 43, 132]. Expenses [128]. Extension [236, 146, 184]. Extent [71, 97]. Extinction [192]. Extract [82]. Extracted [212, 200]. Extracts [10].

F [202, 290, 105, 168, 262, 312, 232, 19]. F.R.S [116]. F.R.S. [116]. Fabric [76].

Factories [323]. Facts [31, 76]. Fallacy [51]. Families [114]. Favour [51]. February [32, 275]. Fellowships [112]. Female [247]. Few [10, 152]. Fields [103]. Fifth [202, 208, 307]. Finances [281, 34]. Financial [89, 150, 190, 222, 251, 278, 307, 47, 119, 122, 127]. First [78, 190, 111]. Fiscal [191]. Florence [273]. following [116]. Force [41]. Foreign [42]. Foreigners [175]. Form [272]. Founded [9, 29]. Fourth [15, 278, 168, 39, 33, 177]. France [54, 228, 252, 275]. Frederick [136]. Free [246]. Freehold [205]. French [55, 133, 143, 216, 225]. Frequent [224]. Friday [82, 251, 47]. Front [110]. Functions [203]. Funds [266]. Future [302, 86].

Gaol [84]. General

[56, 98, 164, 207, 219, 237, 151, 185]. Gentry [264]. Geographical [212, 200]. Germany [12]. Given [264]. Gold [137, 42, 49]. Government [40, 272]. Governors [75]. Great [40]. Greece [285]. Growth [84]. Guy [61].

Hague [321]. Health [67, 291]. Height [84]. Held [82, 89, 138, 150, 267, 278, 296, 307, 47, 273, 321, 39, 23, 78, 106, 116, 177, 190, 208, 222, 239, 251, 182, 33, 119]. Help [40]. Hendriks [136]. Highlands [197, 233]. History [235]. Hosiery [234, 265]. Hospital [61]. Hospitals [98, 164, 207, 219, 237, 253]. House [322]. Household [76]. Human [84].

if [203]. II [212]. Illegitimacy [92]. Illustrated [282, 84]. Illustrative [31]. Impots [48]. Improvements [162]. Improving [20]. Inaccuracies [186]. Inaugural [320]. Incidentally [253]. Inclosure [94]. Inclusive [238]. Income [30, 53, 191, 51, 96, 123]. Inconsistencies [185]. Increase [192]. Increases [286].

India [75, 77, 133, 176, 302, 187, 264, 4]. Indians [291]. Indicating [52]. Indirect [41]. Indiscriminating [51]. Industrial [137]. Infant [186]. Infants [215]. Inferences [29]. Information [1, 28, 38, 65, 81, 101, 109, 131, 142, 167, 181, 201, 213, 231, 243, 261, 271, 289, 300, 319, 10]. Inhabitants [35]. Inquiry [20, 294, 320]. Insanity [308]. Institutions [253, 184, 264]. Instruction [263]. Insufficient [126]. International [295, 15, 24, 20, 135, 147, 273, 321, 39, 33]. Investigation [33]. Ireland [62, 170, 102, 85, 72]. Irregularity [49]. Islands [197]. Issued [238]. Italy [223]. IX [33].

Jame [190, 222]. James [82, 89, 116, 150, 251, 278, 307, 47, 320, 119, 235]. January [238]. Jouissances [48]. Judicial [257]. July [23, 39, 33]. June [23, 19].

Kauyageh [291]. Kingdom [285, 183, 223, 309, 279, 304, 323, 172, 295, 42, 3, 43, 5, 41, 44, 132].

Labour [244]. Labourers [60, 102]. Lace [234]. Lake [145]. Lancashire [2, 97]. Land [205]. Large [206, 218]. Last [66, 61, 68, 121]. Law [258, 170, 111, 16]. Laws [54, 120]. Legislative [275]. Life [136]. Like [133]. List [212, 200, 238]. Live [172]. Liverpool [84]. Lloyd [212, 200, 238]. Local [185]. Locality [133]. London [82, 98, 253, 39, 33, 320, 169]. Lower [91, 245].

M.P [116]. Machinery [266]. Majorities [18]. Male [247]. Males [84]. Manchester [78, 2]. Manner [18]. Manufacture [204, 113, 72]. Manufacturing [2, 247]. Maori [35]. March [190, 222, 278, 307, 89, 150, 190, 222, 251, 278, 307, 47, 238, 119]. Market [224]. Marriage [176]. Marriages [225]. Massachusetts [21]. Material [286]. Matter [110]. Matters [115]. Meaning [203]. Measure [67]. Meeting [23, 8, 78, 82, 89, 106, 116, 135, 138, 150, 147, 177, 202, 190, 208, 222, 239, 251, 267, 278, 296, 307, 47, 290, 105, 182, 168, 262, 312, 119, 232, 19]. Memorandum [52, 32, 41]. Metallic [295]. Method [18]. Methods [115, 33]. Metropolis [253]. Metropolitan [164, 207, 219, 237]. Midland [265]. Military [252]. Mining [137]. Minster [76]. Minutes [82]. Miscellanea [6, 13, 26, 36, 45, 57, 63, 79, 87, 99, 107, 117, 124, 129, 140, 148, 156, 165, 179, 188, 194, 198, 210, 220, 226, 229, 241, 249, 255, 259, 269, 276, 283, 287, 298, 305, 310, 317, 324]. Modern [184]. Modes [41]. Monday [278, 307]. Money [224]. Moral [286]. Mortality [114, 176, 67, 186, 214, 253, 159, 169]. Mr. [74]. Municipal [217]. Muslin [72].

Nassau [19]. Natal [301]. National [95, 182]. Native [264]. Nature [203]. Navies [216]. Navigation [111]. Navy [153]. Neglected [40]. Netherlands [309, 136]. Newcastle [138]. Newcastle-Upon-Tyne [138]. Newmarch [116]. Nine [76]. Nineteenth [114]. Ninth [312, 119]. Norfolk [292]. North [97]. Northumberland [134]. Norwich [296, 290]. Notes [301, 83]. Nottingham [239, 234, 232]. November [320]. Number [175]. Numerical [59, 61].

Observable [49]. Observations [20, 84, 35, 34, 92, 5, 44, 76, 51, 225]. Occupation [175]. Occupations [313]. October [273, 106, 105]. Office [73]. Official [75, 135, 147, 29, 10, 136]. Olmsted [74]. Opening [15, 202, 290, 105, 182, 168, 262, 146, 312, 232, 19]. Operation [54]. operative [71]. Organisation [143, 216]. Organization [20, 303]. Origin [59].

Original [203]. Other [206, 95, 97, 218, 313, 247]. Our [133]. Oxford [23, 235, 112, 19].

Paper [304]. Paris [169, 162]. Parish [134]. Parliament [235]. Part [212, 200]. Particularly [76, 225]. Parts [184, 97, 173]. Passing [16]. Past [128]. Patent [120]. Patients [61]. Pauperism [286, 97, 85]. Pay [123]. Peerage [114]. People [313]. Percentages [238]. Phenomena [49]. Plague [292]. Plea [308]. Poor [170, 16, 16]. Popular [86]. Population [280, 32, 163, 68, 225, 313, 247]. Position [20]. Post [73]. Potables [155]. Powers [94]. Practical [41]. Preliminary [62]. Prepared [33]. Present [20, 31, 128, 34, 294, 86, 22, 41, 320, 74, 83]. Presented [89, 150, 190, 222, 251, 278, 307, 47, 119, 275]. President [15, 116, 202, 290, 105, 168, 262, 312, 232, 19]. Pressure [224, 316]. Prevailing [114, 21]. Prevention [104]. Price [42, 144, 155]. Prices [235, 193, 69, 33, 5, 44, 76]. Prince [15]. Principal [214, 33, 215]. Principally [76]. Principle [94]. Principles [30, 170]. Printing [69]. Prison [91, 245]. Prisoners [126]. Prisons [253]. Procedure [258]. Proceedings [25, 89, 139, 178, 150, 209, 190, 240, 222, 268, 251, 297, 278, 307, 47, 39, 119]. Production [49]. Productiveness [3]. Profits [70]. Programme [33]. Progress [183, 223, 137, 302, 204, 315, 145, 43, 95, 66, 320, 248, 68, 83, 4, 136]. Promotion [182]. Proper [203]. Property [316]. Proportion [247]. Prospective [3]. Prospects [302, 83]. Prosperity [286]. Province [17]. Provincial [164, 207, 219, 237]. Prsident [182]. Prussian [10]. Public [115, 263, 253, 3, 187, 264]. Published [10]. Purposes [132].

Quarterly [7, 14, 27, 37, 46, 58, 64, 80, 88, 93, 100, 108, 118, 125, 130, 141, 149, 157, 166, 180, 189, 195, 199, 211, 221, 227, 230, 242, 250, 256, 260, 270, 277, 284, 288, 299, 306, 311, 318, 325]. Queries [215]. Question [203, 295].

Races [274]. Railway [236, 12]. Raising [41]. Rapid [4]. Rate [114, 2, 16, 123]. Rates [176]. Rationale [120]. Read [32, 116]. Real [316]. Recommendations [11]. Reconsidered [96]. Record [212, 200]. Records [84]. Red [31]. Reference [29, 126, 257, 295, 3, 41, 322, 244]. Reformatories [29]. Regarding [301]. Regards [11, 33]. Regions [74]. Registrar [151, 185]. Registrar-General [151, 185]. Registry [185]. Regular [184]. Relating [257, 152, 293, 41]. Relation [294]. Relative [84, 5, 44, 85]. Remarks [10, 323, 49]. Repeal [54]. Replies [215]. Report [50, 89, 150, 190, 222, 251, 278, 307, 47, 273, 321, 39, 254, 119, 275]. Reported [238]. Reports [75, 135, 147, 151, 185]. Request [33]. Resources [154, 145, 86]. Respecting [10]. Results [62, 56, 236, 170, 71, 5, 44, 151, 29]. Return [37, 323]. Returns [7, 14, 27, 46, 58, 64, 80, 88, 93, 100, 108, 118, 125, 130, 141, 149, 157, 166, 180, 189, 195, 199, 211, 221, 227, 230, 242, 250, 256, 260, 270, 277, 284, 288, 299, 306, 311, 318, 325, 29, 257, 9]. Revenue [3, 41]. Review [9, 32]. Rise [302]. Rochdale [71]. Rogers [235]. Rolls [76]. Roman [158]. Rooms [89, 116, 150, 190, 222, 251, 278, 307, 47, 320, 119]. Royal [153]. Russia [161, 22]. Russian [59].

Salford [2]. Salisbury [174]. Sanitary [133, 67, 174]. Saskatchewan [145]. Saturday [89, 119]. Saving [73]. Scale [54]. Schools [187]. Science [202, 290, 105, 168, 262, 203, 66, 312, 232, 19, 23, 78, 106, 138, 177, 202, 208, 239, 267, 296, 290, 105, 182, 168, 262, 312, 232]. Scotland [170, 197, 102, 85, 72, 68, 233]. Sea [31]. Second [106, 222, 304, 105]. Section [202, 290, 105, 168, 262, 33, 312, 232, 19]. Sections [212, 200]. Senate [275]. September [138, 177, 208, 267, 273, 321, 78, 135, 147, 202, 182, 168, 262]. Serfdom [59, 22]. Service [77, 115, 187, 293, 303]. Session [15, 39, 33]. Sessional [89, 150, 190, 222, 251, 278, 307, 47, 119]. Settlement [291]. Seven [47, 323, 218, 61]. Seventeenth [76]. Seventh [56, 267, 321, 262]. Shipbuilding [314]. Shipping [212, 152, 200, 294]. Shipwrecks [153]. Shown [68]. Shuttleworth [76]. Sickness [67]. Silver [42]. Since [49, 44, 193, 5, 16, 72, 83]. Sixteenth [76]. Sixth [8, 239, 273, 232]. Slave [74]. Sliding [54]. Small [204]. Social [182, 203]. Societies [205]. Society [203, 25, 11, 8, 82, 89, 116, 139, 178, 150, 209, 190, 240, 222, 268, 251, 297, 278, 307, 47, 39, 32, 320, 119]. Soldier [123]. Some [52, 20, 293, 51, 173, 247]. Spain [9, 32]. Spanish [9]. Special [82, 29, 126, 257]. Square [82, 89, 116, 150, 190, 222, 251, 278, 307, 47, 320, 119]. St [116]. St. [82, 89, 150, 190, 222, 251, 278, 307, 47, 320, 119]. State [286, 35, 291, 275]. Statement [49]. Statements [55, 31, 5, 44]. States [280, 214, 10, 215, 281, 21, 74, 122, 127]. Statistical [15, 24, 11, 20, 82, 116, 135, 147, 223, 273, 321, 134, 84, 203, 39, 18, 294, 301, 49, 33, 254, 320, 191, 25, 8, 89, 139, 178, 150, 209, 190, 240, 222, 268, 251, 297, 278, 307, 47, 10, 39, 32, 238, 119]. Statistician [17]. Statistics [52, 98, 164, 202, 207, 219, 237, 282, 285, 133, 258, 290, 309, 206, 279, 304, 105, 168, 111, 262, 203, 257, 9, 90, 32, 172, 152, 12, 92, 158, 293, 161, 174, 91, 245, 33, 312, 16, 232, 218, 19, 264, 160, 122, 127, 225, 313, 171, 136]. Status [175]. Stock [172]. Street [162]. Strength [216]. Strikes [244, 70]. Study [184]. Subject [115]. Subjects [184, 313].

Suez [31]. Sufficient [126]. Suggestions [20]. Summary [56, 238]. Sums [264]. Supplies [42]. Supposed [192]. Sweden [90]. Sykes [116]. System [272, 21, 53, 173]. Systems [41].

Table [136]. Tables [10]. Tabulated

[212, 200]. Tariffs [121]. Tax [30, 53, 51, 96]. Taxation [21, 41, 316, 122, 127]. Taxes [48, 3]. Tea [302]. Ten [121]. Term [203]. Thames [314]. Their [162, 170, 70]. Thence [41]. There [203]. Third [138, 251]. Thirteenth [150]. Thirtieth [23]. Thirty [78, 106, 138, 177, 202, 190, 208, 222, 239, 251, 267, 278, 296, 307, 290, 105, 84, 168, 262, 66, 312, 76, 232]. Thirty-Eighth [296, 290]. Thirty-Fifth [202, 208, 307]. Thirty-First [78, 190]. Thirty-Fourth [278, 168, 177]. Thirty-Nine [76]. Thirty-Ninth [312]. Thirty-Second [106, 222, 105]. Thirty-Seventh [267, 262]. Thirty-Sixth [239, 232]. Thirty-Third [138, 251]. Thorold [235]. Those [133, 9, 225]. Thursday [116, 222]. Time [22]. Tonnage [111]. Topographical [282]. Total [54]. Towns [206, 218, 247]. Trade [31, 274, 323, 113, 5, 44, 83, 246, 4, 121, 182]. Trades [234, 33, 265]. Trading [71]. Traffic [12]. Translator [10]. Treated [61]. Treatment [215]. Troops [133]. Truck [173]. Tuesday [150, 320]. Turkey [192, 263]. Turks [192]. Twenty [8, 89, 47, 119]. Twenty-Eighth [89]. Twenty-Ninth [119]. Twenty-Seven [47]. Twenty-Sixth [8]. Tyne [138].

Undiscriminating [96]. Unequal [247]. Unions [97]. United [183, 279, 304, 281, 323, 172, 21, 295, 42, 3, 43, 5, 41, 44, 74, 132, 122, 127]. University [128, 146, 272]. Upland [69]. Upon [138, 48, 92, 29, 275]. Utility [103, 264].

Value [193]. Variation [193]. Various [52, 121]. Vice [116]. Vice-President [116]. Victoria [137]. Vital [133, 90, 218]. Volume [1, 28, 38, 65, 81, 101, 109, 131, 142, 167, 181, 201, 213, 231, 243, 261, 271, 289, 300, 319]. Votes [18].

W [19]. Wages [2, 33, 76, 123, 244, 70]. Wales [50, 56, 29, 257, 158, 86, 322, 60, 163]. War [113, 235]. Warburton [51]. was [116]. Wednesday [190]. Weight [84]. West [75, 173]. Wheat [144]. Whether [203]. Which [52, 94, 41]. William [116]. Winnipeg [145]. word [191]. Working [120]. Works [264]. Wrecks [238].

Year [89, 150, 190, 212, 235, 222, 251, 278, 307, 47, 10, 200, 5, 44, 119]. Years [29, 84, 323, 9, 174, 66, 76, 144, 61, 121]. York [182, 76].

Zealand [35, 315].

References

Anonymous:1860:VIa

[1] Anonymous. Volume information. Journal of the Statistical Society of London, 23(1):i–viii, March 1860. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338477.

Chadwick:1860:RWM

[2] David Chadwick. On the rate of wages in Manchester and Salford, and the manufacturing districts of Lancashire, 1839– 59. Journal of the Statistical Society of London, 23(1):1–36, March 1860. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338478.

Levi:1860:DPT

[3] Leone Levi. On the distribution and productiveness of taxes with reference to the prospective ameliorations in the public revenue of the United Kingdom. Journal of the Statistical Society of London, 23(1):37–65, March 1860. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338479.

Valpy:1860:RRP

[4] Richard Valpy. On the recent and rapid progress of the British trade with India. Journal of the Statistical Society of London, 23(1):66–75, March 1860. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338480.

Newmarch:1860:RTU

[5] William Newmarch. Results of the trade of the United Kingdom during the year 1859; with statements and observations relative to the course of prices since the year 1844. Journal of the Statistical Society of London, 23(1):76–110, March 1860. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338481.

Anonymous:1860:Ma

[6] Anonymous. Miscellanea. Journal of the Statistical Society of London, 23(1):111– 116, March 1860. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338482.

Anonymous:1860:QRa

[7] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 23(1):117–140, March 1860. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338483.

Anonymous:1860:TSA

[8] Anonymous. Twenty-sixth anniversary meeting of the Statistical Society. Journal of the Statistical Society of London, 23(2):141–146, June 1860. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338564.

Hendricks:1860:RSS

[9] Frederick Hendricks. A review of the statistics of Spain down to the years 1857 and 1858; chie founded on the Spanish census returns of those years. Journal of the Statistical Society of London, 23(2):147–200, June 1860. CODEN ???? ISSN 0959-5341. URL http:/

/www.jstor.org/stable/2338565. See

[32].

Goldsmid:1860:ETO

[10] Francis Henry Goldsmid. Extracts from the tables and official information respecting the Prussian states for the year 1849, published by the Statistical Department at Berlin, and a few remarks by the translator. Journal of the Statistical Society of London, 23(2):201–221, June 1860. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338566.

Anonymous:1860:RCS

[11] Anonymous. Recommendations of the Council of the Statistical Society as regards the census of 1861. Journal of the Statistical Society of London, 23(2):222– 223, June 1860. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338567.

Lazarus:1860:SRE

[12] Wilhelm Lazarus. On the statistics of railway enterprise and traffic in Germany. Journal of the Statistical Society of London, 23(2):224–232, June 1860. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338568.

Anonymous:1860:Mb

[13] Anonymous. Miscellanea. Journal of the Statistical Society of London, 23(2):233– 251, June 1860. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338569.

Anonymous:1860:QRb

[14] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 23(2):253–275, June 1860. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338570.

Anonymous:1860:APC

[15] Anonymous. The address of the Prince Consort on opening as President the fourth session of the International Statistical Congress. Journal of the Statistical Society of London, 23(3):277–285, September 1860. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338498.

Purdy:1860:SEP

[16] Frederick Purdy. The statistics of the English poor rate before and since the passing of the Poor Law Amendment Act. Journal of the Statistical Society of London, 23(3):286–329, September 1860. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338499.

Fox:1860:PS

[17] J. J. Fox. On the province of the statistician. Journal of the Statistical Society of London, 23(3):330–336, September 1860. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338500.

Hare:1860:ANS

[18] Thomas Hare. On the application of a new statistical method to the ascertainment of the votes of majorities in a more exhaustive manner. Journal of the Statistical Society of London, 23 (3):337–356, September 1860. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338501.

Senior:1860:OAN

[19] Nassau W. Senior. Opening address of Nassau W. Senior, Esq., as President of Section F (Economic Science and Statistics), at the meeting of the British Association, at Oxford, 28th June, 1860. Journal of the Statistical Society of London, 23(3):357–361, September 1860. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338502.

Anonymous:1860:SOP

[20] Anonymous. Some observations on the present position of statistical inquiry, with suggestions for improving the organization and efficiency of the International Statistical Congress. Journal of the Statistical Society of London, 23 (3):362–369, September 1860. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338503.

Jarvis:1860:STP

[21] Edward Jarvis. On the system of taxation prevailing in the United States, and especially in Massachusetts. Journal of the Statistical Society of London, 23 (3):370–378, September 1860. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338504.

Michelsen:1860:SRP

[22] Dr. Michelsen. Serfdom in Russia at the present time. Journal of the Statistical Society of London, 23(3):379–381, September 1860. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338505.

Anonymous:1860:TMB

[23] Anonymous. Thirtieth meeting of the British Association for the Advancement of Science, held at Oxford, 27th June– 4th July, 1980. Journal of the Statistical Society of London, 23(3):382–383, September 1860. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338506.

Anonymous:1860:ISC

[24] Anonymous. International Statistical Congress, 1860. Journal of the Statistical Society of London, 23(3):384–386, September 1860. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338507.

Anonymous:1860:PSS

[25] Anonymous. Proceedings of the Statistical Society. Journal of the Statistical Society of London, 23(3):387–389, September 1860. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338508.

Anonymous:1860:Mc

[26] Anonymous. Miscellanea. Journal of the Statistical Society of London, 23(3): 390–403, September 1860. CODEN ????

ISSN 0959-5341. URL http://www.

jstor.org/stable/2338509.

Anonymous:1860:QRc

[27] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 23 (3):404–426, September 1860. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338510.

Anonymous:1860:VIb

[28] Anonymous. Volume information. Journal of the Statistical Society of London, 23(4):579–593, December 1860. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338526.

Baker:1860:AIF

[29] T. Barwick Lloyd Baker. Abstracts and inferences founded upon the official criminal returns of England and Wales for the years 1854–9, with special reference to the results of reformatories. Journal of the Statistical Society of London, 23(4):427–454, December 1860. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338527.

Booth:1860:PIT

[30] Rev. Dr. Booth. On the principles of an income tax. Journal of the Statistical Society of London, 23(4):455–464, December 1860. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338528.

Dassy:1860:FSI

[31] G. F. Dassy. Facts and statements illustrative of the trade of Suez, and of the commerce of the Red Sea, as at present carried on. Journal of the Statistical Society of London, 23(4):465–474, December 1860. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338529.

Hendriks:1860:MPS

[32] Mr. Frederick Hendriks. Memorandum on the population statistics of Spain in 1858 and 1859; in continuation of the “Review of the statistics of Spain, read to the Statistical Society in February, 1860”. Journal of the Statistical Society of London, 23(4):475–478, December 1860. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/

2338530. See [9].

Newmarch:1860:MIR

[33] William Newmarch. On methods of investigation as regards statistics of prices, and of wages in the principal trades. Being the programme prepared by request for the section (IX) commercial statistics, of the fourth session (1860), of the International Statistical Congress held in London in July, 1860. Journal of the Statistical Society of London, 23 (4):479–497, December 1860. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338531.

Horn:1860:OPR

[34] J. E. Horn. Observations on the present and recent condition of the finances of Austria. Journal of the Statistical Society of London, 23(4):498–507, December 1860. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338532.

Fenton:1860:OSA

[35] F. D. Fenton. Observations on the state of the Aboriginal Maori inhabitants of New Zealand. Journal of the Statistical Society of London, 23(4):508–541, December 1860. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338533.

Anonymous:1860:Md

[36] Anonymous. Miscellanea. Journal of the Statistical Society of London, 23(4): 542–555, December 1860. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338534.

Anonymous:1860:QRd

[37] Anonymous. Quarterly return. Journal of the Statistical Society of London, 23 (4):556–578, December 1860. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338535.

Anonymous:1861:VIa

[38] Anonymous. Volume information. Journal of the Statistical Society of London, 24(1):i–viii, March 1861. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338407.

Hammack:1861:RSS

[39] James T. Hammack. Report to the Statistical Society on the proceedings of the fourth session of the International Statistical Congress, held in London, July, 1860. Journal of the Statistical Society of London, 24(1):1–21, March 1861. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338408.

Carpenter:1861:EHG

[40] Miss Carpenter. On educational help from government for the destitute and neglected children of Great Britain. Journal of the Statistical Society of London, 24(1):22–29, March 1861. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338409.

Newmarch:1861:MRS

[41] William Newmarch. Memorandum relating to the systems of taxation at present in force in the United Kingdom, with especial reference to which direct and indirect modes of raising revenue are employed, and the practical effects thence arising. Journal of the Statistical Society of London, 24(1):30–37, March 1861. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338410.

Jourdan:1861:EGS

[42] Francis Jourdan. The effect of the gold supplies on the foreign exchanges between the United Kingdom and foreign countries, and on the price of silver. Journal of the Statistical Society of London, 24(1):38–54, March 1861. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338411.

Levi:1861:PEU

[43] Leone Levi. On the progress of the expenditure of the United Kingdom. Journal of the Statistical Society of London, 24(1):55–73, March 1861. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338412.

Newmarch:1861:RTU

[44] William Newmarch. Results of the trade of the United Kingdom during the year 1860; with statements and observations relative to the course of prices since the year 1844. Journal of the Statistical Society of London, 24(1):74–124, March 1861. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338413.

Anonymous:1861:Ma

[45] Anonymous. Miscellanea. Journal of the Statistical Society of London, 24(1):125– 138, March 1861. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338414.

Anonymous:1861:QRa

[46] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 24(1):139–159, March 1861. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338415.

Boileau:1861:RCF

[47] John P. Boileau. Report of the Council for the financial year ended 31st December, 1860, and for the sessional year ended March, 1861, presented at the twenty-seven anniversary meeting of the Statistical Society, held in the Society’s rooms St. James’s Square, on Friday, 15th March, 1861; — with the proceedings of that meeting. Journal of the Statistical Society of London, 24(2):161– 166, June 1861. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338442.

DeParieu:1861:TUE

[48] M. Esquirou De Parieu. On taxes upon enjoyments (impots sur les jouissances). Journal of the Statistical Society of London, 24(2):167–197, June 1861. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338443.

McArthur:1861:RIS

[49] James McArthur, Edward Hamilton, Stuart Donaldson, W. Marsh, W. Westgarth, E. Stephens, J. A. Youl, and J. E.

Fitzgerald. Remarks on the irregularity of the statistical phenomena observable in the Australian colonies since the gold discovery of 1851; and statement of the production of gold in Australia to the close of 1859. Journal of the Statistical Society of London, 24(2):198– 207, June 1861. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338444.

Anonymous:1861:REE

[50] Anonymous. Report on the education in England and Wales in 1858–60, as ascertained by the Education Commission of 1858. Journal of the Statistical Society of London, 24(2):208–212, June 1861. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338445.

Sargant:1861:SOF

[51] William Lucas Sargant. Some observations on the fallacy of the Warburton argument in favour of an indiscriminating income tax. Journal of the Statistical Society of London, 24(2):213–228, June 1861. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338446.

Alex:1861:MIS

[52] John Alex. Memorandum indicating some of the causes to which the discordancy of the commercial statistics of various countries in chiefly to be attributed. Journal of the Statistical Society of London, 24(2):229–234, June 1861. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338447.

Michelsen:1861:DSI

[53] Dr. Michelsen. The decimal system of the income tax and the budgets of the ancients. Journal of the Statistical Society of London, 24(2):235–236, June 1861. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338448.

Anonymous:1861:OSS

[54] Anonymous. Operation of the sliding scale of corn laws in France, and its approaching total repeal. Journal of the Statistical Society of London, 24(2):237– 240, June 1861. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338449.

Anonymous:1861:BFA

[55] Anonymous. The British and French armies. Comparative statements, 1860– 61. Journal of the Statistical Society of London, 24(2):241–246, June 1861. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338450.

Anonymous:1861:SCE

[56] Anonymous. Seventh census of England and Wales, April, 1861. summary of general results. Journal of the Statistical Society of London, 24(2):247–272, June 1861. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338451.

Anonymous:1861:Mb

[57] Anonymous. Miscellanea. Journal of the Statistical Society of London, 24(2):273– 287, June 1861. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338452.

Anonymous:1861:QRb

[58] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 24(2):288–309, June 1861. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338453.

DeBuschen:1861:OND

[59] M. Arthur De Buschen. On the origin and numerical development of serfdom in the Russian empire. Journal of the Statistical Society of London, 24 (3):311–327, September 1861. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338484.

Purdy:1861:EAL

[60] Frederick Purdy. On the earnings of agricultural labourers in England and Wales, 1860. Journal of the Statistical Society of London, 24(3):328–373, September 1861. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338485.

Steele:1861:NAP

[61] John Charles Steele. Numerical analysis of the patients treated in Guy’s hospital for the last seven years, from 1854 to 1861. Journal of the Statistical Society of London, 24(3):374–401, September 1861. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338486.

Anonymous:1861:CIA

[62] Anonymous. Census of Ireland, April, 1861.-preliminary results. Journal of the Statistical Society of London, 24(3):402– 410, September 1861. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338487.

Anonymous:1861:Mc

[63] Anonymous. Miscellanea. Journal of the Statistical Society of London, 24(3): 411–430, September 1861. CODEN ????

ISSN 0959-5341. URL http://www.

jstor.org/stable/2338488.

Anonymous:1861:QRc

[64] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 24 (3):431–450, September 1861. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338489.

Anonymous:1861:VIb

[65] Anonymous. Volume information. Journal of the Statistical Society of London, 24(4):637–654, December 1861. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338461.

Newmarch:1861:PES

[66] William Newmarch. The progress of economic science during the last thirty years. Journal of the Statistical Society of London, 24(4):451–471, December 1861. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338462.

Farr:1861:HBA

[67] Dr. Farr. The health of the British Army, and the effects of recent sanitary measure on its mortality and sickness. Journal of the Statistical Society of London, 24(4):472–484, December 1861. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338463.

Strang:1861:CPP

[68] John Strang. On the comparative progress of the population of England and Scotland, as shown by the last (1861) census. Journal of the Statistical Society of London, 24(4):485–490, December 1861. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338464.

Neild:1861:APP

[69] Alderman Neild. An account of the prices of printing cloth and upland cotton, from 1812 to 1860. Journal of the Statistical Society of London, 24(4): 491–497, December 1861. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338465.

Watts:1861:STE

[70] John Watts. On strikes and their effects on wages, profits, and accumulations. Journal of the Statistical Society of London, 24(4):498–506, December 1861. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338466.

Molesworth:1861:ERC

[71] W. N. Molesworth. On the extent and results of co-operative trading associations at Rochdale. Journal of the Statistical Society of London, 24(4):507–514, December 1861. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338467.

Strang:1861:ACE

[72] John Strang. On the altered condition of the embroidered muslin manufacture of Scotland and Ireland since 1857. Journal of the Statistical Society of London, 24 (4):515–518, December 1861. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338468.

Chadwick:1861:POS

[73] Edwin Chadwick. Post Office savings’ banks. Journal of the Statistical Society of London, 24(4):519–522, December 1861. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338469.

Olmsted:1861:MOA

[74] Mr. Olmsted. Mr. Olmsted’s account of the present condition of the slave regions of the United States. Journal of the Statistical Society of London, 24 (4):523–527, December 1861. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338470.

Anonymous:1861:BWI

[75] Anonymous. British West India colonies in 1859. abstract of the official reports by the governors. Journal of the Statistical Society of London, 24(4):528–534, December 1861. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338471.

Rogers:1861:FOW

[76] James E. T. Rogers. Facts and observations on wages and prices in England during the Sixteenth and Seventeenth Centuries, and more particularly during the thirty-nine years 1582–1620; the date principally employed being the fabric rolls of York Minster and the Shuttleworth household books. Journal of the Statistical Society of London, 24 (4):535–585, December 1861. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338472.

Anonymous:1861:ECA

[77] Anonymous. Effects of competitive appointments in the civil service of India. Journal of the Statistical Society of London, 24(4):586–589, December 1861. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338473.

Anonymous:1861:TFM

[78] Anonymous. Thirty-first meeting of the British Association for the Advancement of Science, held at Manchester, 4th–11th September, 1861. Journal of the Statistical Society of London, 24(4):590–592, December 1861. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338474.

Anonymous:1861:Md

[79] Anonymous. Miscellanea. Journal of the Statistical Society of London, 24(4): 593–615, December 1861. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338475.

Anonymous:1861:QRd

[80] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 24 (4):616–636, December 1861. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338476.

Anonymous:1862:VIa

[81] Anonymous. Volume information. Journal of the Statistical Society of London, 25(1):i–viii, March 1862. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338556.

Anonymous:1862:EMS

[82] Anonymous. Extract from the minutes of a special meeting of the Council of the Statistical Society, held at 12, St. James’s Square, London, on Friday, the 20th December, 1861. Journal of the Statistical Society of London, 25(1):1– 2, March 1862. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338557.

Sykes:1862:NPT

[83] Colonel Sykes. Notes on the progress of the trade of England with China since 1833, and on its present condition and prospects. Journal of the Statistical Society of London, 25(1):3–19, March 1862. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338558.

Danson:1862:SOR

[84] J. T. Danson. Statistical observations relative to the growth of the human body (males) in height and weight, from eighteen to thirty years of age, as illustrated by the records of the Borough Gaol of Liverpool. Journal of the Statistical Society of London, 25(1):20–26, March 1862. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338559.

Purdy:1862:RPE

[85] Frederick Purdy. The relative pauperism of England, Scotland, and Ireland, 1851 to 1860. Journal of the Statistical Society of London, 25(1):27–49, March 1862. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338560.

Mann:1862:RPE

[86] Horace Mann. The resources of popular education in England and Wales: Present and future. Journal of the Statistical Society of London, 25(1):50–71, March 1862. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338561.

Anonymous:1862:Ma

[87] Anonymous. Miscellanea. Journal of the Statistical Society of London, 25(1):72– 83, March 1862. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338562.

Anonymous:1862:QRa

[88] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 25(1):84–104, March 1862. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338563.

Anonymous:1862:RCF

[89] Anonymous. Report of the Council for the financial year ended 31st December, 1861, and for the sessional year ended March 1862, presented at the twentyeighth anniversary meeting of the Statistical Society, held at the Society’s rooms, 12, St. James’s Square, on Saturday, 15th March, 1862; with the proceedings of that meeting. Journal of the Statistical Society of London, 25(2):105– 110, June 1862. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338402.

Hendricks:1862:VSS

[90] Frederick Hendricks. On the vital statistics of Sweden, from 1749 to 1855. Journal of the Statistical Society of London, 25(2):111–174, June 1862. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338403.

Mouat:1862:PSD

[91] Frederic John Mouat. On prison statistics and discipline in lower Bengal. Journal of the Statistical Society of London, 25(2):175–218, June 1862. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338404.

Lumley:1862:OUS

[92] W. G. Lumley. Observations upon the statistics of illegitimacy. Journal of the Statistical Society of London, 25(2):219– 274, June 1862. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338405.

Anonymous:1862:QRb

[93] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 25(2):275–296, June 1862. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338406.

Tottie:1862:ICP

[94] John William Tottie. On the Inclosure Commission, its powers, and the principle on which they have been exercised. Journal of the Statistical Society of London, 25(3):297–312, September 1862. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338536.

Levi:1862:PEB

[95] Leone Levi. On the progress and economical bearings of national debts in this and other countries. Journal of the Statistical Society of London, 25(3): 313–338, September 1862. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338537.

Sargant:1862:UIT

[96] William Lucas Sargant. An undiscriminating income tax reconsidered. Journal of the Statistical Society of London, 25 (3):339–376, September 1862. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338538.

Purdy:1862:EPD

[97] Frederick Purdy. Extent of pauperism in the distressed unions in Lancashire and other parts of the North of England, 1861–62. Journal of the Statistical Society of London, 25(3):377–383, September 1862. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338539.

Anonymous:1862:SGH

[98] Anonymous. Statistics of the general hospitals of London, 1861. Journal of the Statistical Society of London, 25(3): 384–388, September 1862. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338540.

Anonymous:1862:Mb

[99] Anonymous. Miscellanea. Journal of the Statistical Society of London, 25(3): 389–404, September 1862. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338541.

Anonymous:1862:QRc

[100] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 25 (3):405–424, September 1862. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338542.

Anonymous:1862:VIb

[101] Anonymous. Volume information. Journal of the Statistical Society of London, 25(4):561–574, December 1862. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338597.

Purdy:1862:EAL

[102] Frederick Purdy. On the earnings of agricultural labourers in Scotland and Ireland. Journal of the Statistical Society of London, 25(4):425–490, December 1862. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338598.

Merivale:1862:UCF

[103] Herman Merivale. On the utility of colonies as fields for emigration. Journal of the Statistical Society of London, 25(4):491–496, December 1862. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338599.

Hill:1862:PC

[104] Edwin Hill. On the prevention of crime. Journal of the Statistical Society of London, 25(4):497–501, December 1862. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338600.

Chadwick:1862:OAP

[105] Edwin Chadwick. Opening address of the President of Section F (Economic Science and Statistics) of the British Association for the Advancement of Science, a the thirty-second meeting, at Cambridge, in October, 1862. Journal of the Statistical Society of London, 25 (4):502–524, December 1862. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338601.

Anonymous:1862:TSM

[106] Anonymous. Thirty-second meeting of the British Association for the Advancement of Science, held at Cambridge, 1st– 8th October, 1862. Journal of the Statistical Society of London, 25(4):525–526, December 1862. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338602.

Anonymous:1862:Mc

[107] Anonymous. Miscellanea. Journal of the Statistical Society of London, 25(4): 527–538, December 1862. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338603.

Anonymous:1862:QRd

[108] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 25 (4):539–559, December 1862. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338604.

Anonymous:1863:VIa

[109] Anonymous. Volume information. Journal of the Statistical Society of London, 26(1):i–viii, March 1863. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338634.

Anonymous:1863:FM

[110] Anonymous. Front matter. Journal of the Statistical Society of London, 26(1): ii, March 1863. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338635.

Glover:1863:STD

[111] John Glover. On the statistics of tonnage during the first decade under the navigation law of 1849. Journal of the Statistical Society of London, 26(1):1– 18, March 1863. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338636.

Heywood:1863:EEO

[112] James Heywood. Endowed education, and Oxford and Cambridge College fellowships. Journal of the Statistical Society of London, 26(1):19–25, March 1863. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338637.

Levi:1863:CTM

[113] Leone Levi. On the cotton trade and manufacture, as affected by the civil war in America. Journal of the Statistical Society of London, 26(1):26–48, March 1863. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338638.

Bailey:1863:RMP

[114] Arthur Hutchinson Bailey and Archibald Day. On the rate of mortality prevailing amongst the families of the peerage during the Nineteenth Century. Journal of the Statistical Society of London, 26(1): 49–71, March 1863. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338639.

Chadwick:1863:SMM

[115] Edwin Chadwick. On the subject matters and methods of competitive examinations for the public service. Journal of the Statistical Society of London, 26(1): 72–77, March 1863. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338640.

Anonymous:1863:MCS

[116] Anonymous. At a meeting of the Council of the Statistical Society, held at the rooms of the Society, 12 St. James’s Square, on Thursday, 11th December, 1862, Colonel Sykes, M.P., F.R.S., Vice-President, in the Chair the following communication from William Newmarch, Esq., F.R.S., was read. Journal of the Statistical Society of London, 26(1):78–81, March 1863. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338641.

Anonymous:1863:Ma

[117] Anonymous. Miscellanea. Journal of the Statistical Society of London, 26(1):82– 90, March 1863. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338642.

Anonymous:1863:QRa

[118] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 26(1):91–114, March 1863. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338643.

Pakington:1863:RCF

[119] John S. Pakington. Report of the Council for the financial year ended 31st December, 1862, and for the sessional year ended March, 1863, presented at the twenty-ninth anniversary meeting of the Statistical Society, held at the Society’s rooms, 12, St. James’s Square, on Saturday, 14th March, 1863; with the proceedings of that meeting. Journal of the Statistical Society of London, 26(2):115– 120, June 1863. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338454.

Rogers:1863:RWP

[120] J. E. T. Rogers. On the rationale and working of the patent laws. Journal of the Statistical Society of London, 26(2): 121–142, June 1863. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338455.

Valpy:1863:TTV

[121] Richard Valpy. The tariffs and trade of various countries during the last ten years. Journal of the Statistical Society of London, 26(2):143–153, June 1863. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338456.

Walford:1863:RFTa

[122] Cornelius Walford. Recent financial and taxation statistics of the United States. Journal of the Statistical Society of London, 26(2):154–167, June 1863. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338457.

Tulloch:1863:PIB

[123] Alexander Murray Tulloch. On the pay and income of the British soldier, as compared with the rate of agricultural wages. Journal of the Statistical Society of London, 26(2):168–185, June 1863. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338458.

Anonymous:1863:Mb

[124] Anonymous. Miscellanea. Journal of the Statistical Society of London, 26(2):186– 211, June 1863. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338459.

Anonymous:1863:QRb

[125] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 26(2):212–237, June 1863. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338460.

Guy:1863:SID

[126] William A. Guy. On sufficient and insufficient dietaries, with special reference to the dietaries of prisoners. Journal of the Statistical Society of London, 26 (3):239–280, September 1863. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338388.

Walford:1863:RFTb

[127] Cornelius Walford. Recent financial and taxation statistics of the United States. Journal of the Statistical Society of London, 26(3):281–295, September 1863. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338389.

Emery:1863:EUE

[128] William Emery. Expenses of university education at Cambridge; past and present. Journal of the Statistical Society of London, 26(3):296–316, September 1863. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338390.

Anonymous:1863:Mc

[129] Anonymous. Miscellanea. Journal of the Statistical Society of London, 26(3): 317–332, September 1863. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338391.

Anonymous:1863:QRc

[130] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 26 (3):333–358, September 1863. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338392.

Anonymous:1863:VIb

[131] Anonymous. Volume information. Journal of the Statistical Society of London, 26(4):487–500, December 1863. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338431.

Purdy:1863:EUK

[132] Frederick Purdy. The expenditure of the United Kingdom for colonial purposes. Journal of the Statistical Society of London, 26(4):359–383, December 1863. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338432.

Bird:1863:VSS

[133] James Bird. On the vital and sanitary statistics of our European Army in India, compared with those of French troops under like conditions of climate and locality. Journal of the Statistical Society of London, 26(4):384–405, December 1863. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338433.

Charlton:1863:SAP

[134] William Henry Charlton. A statistical account of the Parish of Bellingham, Northumberland. Journal of the Statistical Society of London, 26(4):406–411, December 1863. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338434.

Anonymous:1863:ROD

[135] Anonymous. Reports of the official delegates from England at the meeting of the International Statistical Congress, Berlin, September, 1863. Journal of the Statistical Society of London, 26(4): 412–419, December 1863. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338435.

vonBaumhauer:1863:POS

[136] Dr. von Baumhauer and Frederick Hendriks. On the progress of official statistics in the Netherlands (1858–63); — with a new Dutch life table by Dr. von Baumhauer; contributed by Frederick Hendriks. Journal of the Statistical Society of London, 26(4):420–423, December 1863. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338436.

Chapman:1863:IPV

[137] H. S. Chapman. The industrial progress of Victoria as connected with its gold mining. Journal of the Statistical Society of London, 26(4):424–442, December 1863. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338437.

Anonymous:1863:TTM

[138] Anonymous. Thirty-third meeting of the British Association for the Advancement of Science, held at Newcastle-Upon-Tyne, 26th August–2nd September, 1863. Journal of the Statistical Society of London, 26(4):443–444, December 1863. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338438.

Anonymous:1863:PSS

[139] Anonymous. Proceedings of the Statistical Society. Journal of the Statistical Society of London, 26(4):445–450, December 1863. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338439.

Anonymous:1863:Md

[140] Anonymous. Miscellanea. Journal of the Statistical Society of London, 26(4): 451–459, December 1863. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338440.

Anonymous:1863:QRd

[141] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 26 (4):460–485, December 1863. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338441.

Anonymous:1864:VIa

[142] Anonymous. Volume information. Journal of the Statistical Society of London, 27(1):i–viii, March 1864. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338589.

Sykes:1864:COC

[143] W. H. Sykes. Comparison of the organisation and cost in detail of the English and French armies. Journal of the Statistical Society of London, 27(1):1– 69, March 1864. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338590.

Rogers:1864:CPW

[144] J. E. T. Rogers. On a continuous price of wheat for 105 years, from 1380 to 1484. Journal of the Statistical Society of London, 27(1):70–81, March 1864. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338591.

Hind:1864:CPR

[145] Henry Youle Hind. On the commercial progress and resources of central British America; the Lake Winnipeg and Saskatchewan districts. Journal of the Statistical Society of London, 27(1):82– 105, March 1864. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338592.

Heywood:1864:OED

[146] James Heywood. On the opening and extension of Durham University academical endowments. Journal of the Statistical Society of London, 27(1):106–111, March 1864. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338593.

Anonymous:1864:ROD

[147] Anonymous. Reports of the official delegates from England at the meeting of the International Statistical Congress, Berlin, September, 1863. Journal of the Statistical Society of London, 27(1):112– 117, March 1864. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338594.

Anonymous:1864:Ma

[148] Anonymous. Miscellanea. Journal of the Statistical Society of London, 27(1):118– 136, March 1864. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338595.

Anonymous:1864:QRa

[149] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 27(1):137–164, March 1864. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338596.

Anonymous:1864:RCF

[150] Anonymous. Report of the Council for the financial year ended 31st December, 1863, and for the sessional year ended March, 1864, presented at the thirteenth anniversary meeting of the Statistical Society, held at the Society’s rooms, 12, St. James’s Square, on Tuesday, 15th March, 1864; with the proceedings of that meeting. Journal of the Statistical Society of London, 27(2):165–169, June 1864. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338548.

Sargant:1864:CRD

[151] William Lucas Sargant. On certain results and defects of the reports of the Registrar-General. Journal of the Statistical Society of London, 27(2):170–221, June 1864. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338549.

Jeula:1864:FSR

[152] Henry Jeula. A few statistics relating to shipping casualities. Journal of the Statistical Society of London, 27(2):222– 233, June 1864. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338550.

Hodge:1864:SRN

[153] William Barwick Hodge. On shipwrecks in the Royal Navy. Journal of the Statistical Society of London, 27(2):234–244, June 1864. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338551.

Heywood:1864:RB

[154] James Heywood. Resources of Brazil.

Journal of the Statistical Society of London, 27(2):245–257, June 1864. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338552.

Sykes:1864:PEP

[155] W. H. Sykes. Price of edibles and potables in A.D. 1506. Journal of the Statistical Society of London, 27(2):258–260, June 1864. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338553.

Anonymous:1864:Mb

[156] Anonymous. Miscellanea. Journal of the Statistical Society of London, 27(2):261– 276, June 1864. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338554.

Anonymous:1864:QRb

[157] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 27(2):277–301, June 1864. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338555.

Lumley:1864:SRC

[158] William Golden Lumley. The statistics of the Roman Catholics in England and Wales. Journal of the Statistical Society of London, 27(3):303–323, September 1864. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338517.

Tait:1864:ME

[159] P. M. Tait. On the mortality of Eurasians. Journal of the Statistical Society of London, 27(3):324–356, September 1864. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338518.

Valentine:1864:SCA

[160] James Valentine and W. H. Sykes. Statistics of the city of Aberdeen. Journal of the Statistical Society of London, 27(3):357–368, September 1864. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338519.

Michell:1864:SCR

[161] T. Michell. Statistics of crime in Russia. Journal of the Statistical Society of London, 27(3):369–377, September 1864. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338520.

Tite:1864:PSI

[162] W. Tite. On the Paris street improvements, and their cost. Journal of the Statistical Society of London, 27(3):378– 387, September 1864. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338521.

Purdy:1864:DAP

[163] Frederick Purdy. On the decrease of the agricultural population of England and Wales, 1851–61. Journal of the Statistical Society of London, 27(3):388–400, September 1864. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338522.

Anonymous:1864:SMP

[164] Anonymous. Statistics of metropolitan and provincial general hospitals for 1862. Journal of the Statistical Society of London, 27(3):401–409, September 1864. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338523.

Anonymous:1864:Mc

[165] Anonymous. Miscellanea. Journal of the Statistical Society of London, 27(3): 410–434, September 1864. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338524.

Anonymous:1864:QRc

[166] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 27 (3):435–458, September 1864. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338525.

Anonymous:1864:VIb

[167] Anonymous. Volume information. Journal of the Statistical Society of London, 27(4):633–648, December 1864. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338608.

Farr:1864:OAP

[168] William Farr. Opening address of the President of Section F (Economic Science and Statistics), of the British Association for the Advancement of Science, at the thirty-fourth meeting, at Bath, in September, 1864. Journal of the Statistical Society of London, 27(4):459–478, December 1864. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338609.

Tite:1864:CML

[169] William Tite. On the comparative mortality of London and Paris. Journal of the Statistical Society of London, 27 (4):479–491, December 1864. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338610.

Chadwick:1864:PLA

[170] Edwin Chadwick. Poor law administration, its Chief principles and their results in England and Ireland as compared with Scotland. Journal of the Statistical Society of London, 27(4):492–504, December 1864. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338611.

Westgarth:1864:SCA

[171] William Westgarth. The statistics of crime in Australia. Journal of the Statistical Society of London, 27(4):505–519, December 1864. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338612.

Herbert:1864:SLS

[172] Robert Herbert. Statistics of live stock in the United Kingdom, 1853–63. Journal of the Statistical Society of London, 27(4):520–525, December 1864. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338613.

Spender:1864:CTS

[173] Edward Spender. On the “cider-truck”

system in some parts of the West of England. Journal of the Statistical Society of London, 27(4):526–532, December 1864. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338614.

Middleton:1864:SSS

[174] A. B. Middleton. The sanitary statistics of Salisbury. The years 1841–49 and 1856–64. Journal of the Statistical Society of London, 27(4):533–557, December 1864. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338615.

Levi:1864:NOS

[175] Leone Levi. On the number, occupation, and status of foreigners in England. Journal of the Statistical Society of London, 27(4):558–565, December 1864. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338616.

Brown:1864:RMM

[176] Samuel Brown. On the rates of mortality and marriage amongst Europeans in India. Journal of the Statistical Society of London, 27(4):566–579, December 1864. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338617.

Anonymous:1864:TFM

[177] Anonymous. Thirty-fourth meeting of the British Association for the Advancement of Science, held at Bath, 15th–20th September, 1864. Journal of the Statistical Society of London, 27(4):580–581, December 1864. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338618.

Anonymous:1864:PSS

[178] Anonymous. Proceedings of the Statistical Society. Journal of the Statistical Society of London, 27(4):582–583, December 1864. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338619.

Anonymous:1864:Md

[179] Anonymous. Miscellanea. Journal of the Statistical Society of London, 27(4): 584–602, December 1864. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338620.

Anonymous:1864:QRd

[180] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 27 (4):603–631, December 1864. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338621.

Anonymous:1865:VIa

[181] Anonymous. Volume information. Journal of the Statistical Society of London, 28(1):i–viii, March 1865. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338393.

Chadwick:1865:OAP

[182] Edwin Chadwick. Opening address of the Prsident of the Department of Economy and Trade, at the meeting of the National Association for the Promotion of Social Science, held at York, in September, 1864. Journal of the Statistical Society of London, 28(1):1–33, March 1865. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338394.

Blakely:1865:CPC

[183] E. T. Blakely. On the commercial progress of the Colonies and Dependencies of the United Kingdom. Journal of the Statistical Society of London, 28(1): 34–55, March 1865. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338395.

Heywood:1865:EMS

[184] James Heywood. On the extension of modern subjects, as parts of regular study, in educational institutions. Journal of the Statistical Society of London, 28(1):56–72, March 1865. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338396.

Sargant:1865:IEC

[185] William Lucas Sargant. Inconsistencies of the English census of 1861, with the Registrar-General’s reports: and deficiencies in the local registry of births.

Journal of the Statistical Society of London, 28(1):73–124, March 1865. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338397.

Farr:1865:IMA

[186] William Farr. On infant mortality, and on alleged inaccuracies of the census. Journal of the Statistical Society of London, 28(1):125–149, March 1865. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338398.

Mann:1865:PSC

[187] Horace Mann. On public schools and the civil service of India. Journal of the Statistical Society of London, 28(1):150– 169, March 1865. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338399.

Anonymous:1865:Ma

[188] Anonymous. Miscellanea. Journal of the Statistical Society of London, 28(1):170– 198, March 1865. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338400.

Anonymous:1865:QRa

[189] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 28(1):199–230, March 1865. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338401.

Anonymous:1865:RCF

[190] Anonymous. Report of the Council for the financial year ended 31st December, 1864, and for the sessional year ended 15th March, 1865, presented at the thirty-first anniversary meeting of the Statistical Society, held at the Society’s rooms, 12, St. James’ Square, on Wednesday, 15th March, 1865; with the proceedings of that meeting. Journal of the Statistical Society of London, 28(2): 231–241, June 1865. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338416.

Rogers:1865:SFD

[191] J. E. Thorold Rogers. On the statistical and fiscal definitions of the word “income”. Journal of the Statistical Society of London, 28(2):242–260, June 1865. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338417.

Clarke:1865:SET

[192] Hyde Clarke. On the supposed extinction of the Turks and increase of the Christians in Turkey. Journal of the Statistical Society of London, 28(2):261– 293, June 1865. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338418.

Jevons:1865:VPV

[193] W. Stanley Jevons. On the variation of prices and the value of the currency since 1782. Journal of the Statistical Society of London, 28(2):294–320 + 1– 4, June 1865. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338419.

Anonymous:1865:Mb

[194] Anonymous. Miscellanea. Journal of the Statistical Society of London, 28(2):321– 332, June 1865. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338420.

Anonymous:1865:QRb

[195] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 28(2):333–359, June 1865. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338421.

Lubbock:1865:CC

[196] John Lubbock. On the “country clearing”. Journal of the Statistical Society of London, 28(3):361–371, September 1865. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338543.

Levi:1865:ECH

[197] Leone Levi. On the economic condition of the highlands and islands of Scotland. Journal of the Statistical Society of London, 28(3):372–401, September 1865. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338544.

Anonymous:1865:Mc

[198] Anonymous. Miscellanea. Journal of the Statistical Society of London, 28(3): 402–423, September 1865. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338545.

Anonymous:1865:QRc

[199] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 28 (3):424–458, September 1865. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338546.

Jeula:1865:SCA

[200] Henry Jeula. Shipping casualities: An approximate record, extracted from “Lloyd’s List” for the year 1864, tabulated and arranged in geographical sections, Part I. Journal of the Statistical Society of London, 28(3):459–470, September 1865. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338547.

Anonymous:1865:VIb

[201] Anonymous. Volume information. Journal of the Statistical Society of London, 28(4):609–622, December 1865. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338571.

Anonymous:1865:OAP

[202] Anonymous. Opening address of the President of Section F (Economic Science and Statistics), of the British Association for the Advancement of Science, at the thirty-fifth meeting, at Birmingham, September, 1865. Journal of the Statistical Society of London, 28(4):471– 477, December 1865. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338572.

Guy:1865:OAM

[203] William A. Guy. On the original and acquired meaning of the term “statistics,” and on the proper functions of a statistical society: also on the question whether there be a science of statistics; an, if so, what are its nature and “social science”. Journal of the Statistical Society of London, 28(4):478–493, December 1865. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338573.

Goodman:1865:PSA

[204] J. D. Goodman. On the progress of the small arms manufacture. Journal of the Statistical Society of London, 28 (4):494–506, December 1865. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338574.

Johnson:1865:BBF

[205] G. J. Johnson. On the Benefit Building and Freehold Land Societies of Birmingham. Journal of the Statistical Society of London, 28(4):507–517, December 1865. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338575.

Bunce:1865:SCB

[206] J. Thackray Bunce. On the statistics of crime in Birmingham, as compared with other large towns. Journal of the Statistical Society of London, 28(4):518–526, December 1865. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338576.

Anonymous:1865:SMP

[207] Anonymous. Statistics of metropolitan and provincial general hospitals for 1863. Journal of the Statistical Society of London, 28(4):527–535, December 1865. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338577.

Anonymous:1865:TFM

[208] Anonymous. Thirty-fifth meeting of the British Association for the Advancement of Science, held at Birmingham, 6th– 13th September, 1865. Journal of the Statistical Society of London, 28(4):536– 537, December 1865. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338578.

Anonymous:1865:PSS

[209] Anonymous. Proceedings of the Statistical Society. Journal of the Statistical Society of London, 28(4):538–539, December 1865. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338579.

Anonymous:1865:Md

[210] Anonymous. Miscellanea. Journal of the Statistical Society of London, 28(4): 540–555, December 1865. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338580.

Anonymous:1865:QRd

[211] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 28 (4):556–587, December 1865. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338581.

Anonymous:1865:SCA

[212] Anonymous. Shipping casualties: An approximate record, extracted from “Lloyd’s List” for the year 1864, tabulated and arranged in geographical sections, Part II. Journal of the Statistical Society of London, 28(4):588–607, December 1865. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338582.

Anonymous:1866:VIa

[213] Anonymous. Volume information. Journal of the Statistical Society of London, 29(1):i–viii, March 1866. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338422.

Farr:1866:MCP

[214] William Farr. Mortality of children in the principal states of Europe. Journal of the Statistical Society of London, 29(1):1–12, March 1866. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338423.

Stark:1866:RQT

[215] Dr. Stark. Replies to queries as to the treatment of infants in the principal states of Europe. Journal of the Statistical Society of London, 29(1):13– 35, March 1866. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338424.

Sykes:1866:OSC

[216] W. H. Sykes. Organisation, strength, and cost of the English and French Navies in 1865. Journal of the Statistical Society of London, 29(1):36–77, March 1866. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338425.

Avery:1866:MEB

[217] Thomas Avery. On the municipal expenditure of the Borough of Birmingham. Journal of the Statistical Society of London, 29(1):78–91, March 1866. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338426.

Sargant:1866:VSB

[218] William Lucas Sargant. On the vital statistics of Birmingham and seven other large towns. Journal of the Statistical Society of London, 29(1):92–111, March 1866. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338427.

Anonymous:1866:SMPa

[219] Anonymous. Statistics of metropolitan and provincial general hospitals for 1864. Journal of the Statistical Society of London, 29(1):112–121, March 1866. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338428.

Anonymous:1866:Ma

[220] Anonymous. Miscellanea. Journal of the Statistical Society of London, 29(1):122– 160, March 1866. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338429.

Anonymous:1866:QRa

[221] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 29(1):161–190, March 1866. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338430.

Anonymous:1866:RCF

[222] Anonymous. Report of the Council for the financial year ended 31st December, 1865, and for the sessional year ended 15th March, 1866, presented at the thirty-second anniversary meeting of the Statistical Society, held at the Society’s rooms, 12, St. James’ Square, on Thursday, 15th March, 1866; with the proceedings of that meeting. Journal of the Statistical Society of London, 29(2): 191–196, June 1866. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338583.

Brown:1866:SPK

[223] Samuel Brown. On the statistical progress of the Kingdom of Italy. Journal of the Statistical Society of London, 29(2):197–234, June 1866. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338584.

Jevons:1866:FAP

[224] W. Stanley Jevons. On the frequent autumnal pressure in the money market, and the action of the Bank of England. Journal of the Statistical Society of London, 29(2):235–253, June 1866. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338585.

Welton:1866:OFP

[225] T. A. Welton. Observations on French population statistics, particularly those of births, deaths, and marriages. Journal of the Statistical Society of London, 29(2):254–281, June 1866. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338586.

Anonymous:1866:Mb

[226] Anonymous. Miscellanea. Journal of the Statistical Society of London, 29(2):282– 290, June 1866. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338587.

Anonymous:1866:QRb

[227] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 29(2):291–322, June 1866. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338588.

Balfour:1866:BAE

[228] Major-General Balfour. On the budgets and accounts of England and France. Journal of the Statistical Society of London, 29(3):323–444, September 1866. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338605.

Anonymous:1866:Mc

[229] Anonymous. Miscellanea. Journal of the Statistical Society of London, 29(3): 445–460, September 1866. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338606.

Anonymous:1866:QRc

[230] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 29 (3):461–492, September 1866. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338607.

Anonymous:1866:VIb

[231] Anonymous. Volume information. Journal of the Statistical Society of London, 29(4):665–679, December 1866. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338644.

Rogers:1866:OAP

[232] James E. Thorold Rogers. Opening address of the President of Section F (Economic Science and Statistics), of the British Association for the Advancement of Science, at the thirty-sixth meeting, at Nottingham, August, 1866. Journal of the Statistical Society of London, 29 (4):493–503, December 1866. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338645.

ofArgyll:1866:ECH

[233] Duke of Argyll. On the economic condition of the highlands of Scotland. Journal of the Statistical Society of London, 29(4):504–535, December 1866. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338646.

Felkin:1866:LHT

[234] William Felkin. The lace and hosiery trades of Nottingham. Journal of the Statistical Society of London, 29(4):536– 541, December 1866. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338647.

Anonymous:1866:HAP

[235] Anonymous. A history of agriculture and prices in England, from the year after the Oxford Parliament, 1259, to the commencement of the Continental War, 1793, by James E. Thorold Rogers. Journal of the Statistical Society of London, 29(4):542–548, December 1866. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338648.

Baxter:1866:RER

[236] R. Dudley Baxter. Railway extension and its results. Journal of the Statistical Society of London, 29(4):549–595, December 1866. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338649.

Anonymous:1866:SMPb

[237] Anonymous. Statistics of metropolitan and provincial general hospitals for 1865. Journal of the Statistical Society of London, 29(4):596–605, December 1866. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338650.

Jeula:1866:APS

[238] Henry Jeula. An analysis with percentages, of the “Summary of wrecks and casualties reported in ‘Lloyd’s List’ from January to March inclusive 1866” issued by the Statistical Committee of Lloyd’s. Journal of the Statistical Society of London, 29(4):606–609, December 1866. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338651.

Anonymous:1866:TSM

[239] Anonymous. Thirty-sixth meeting of the British Association for the Advancement of Science, held at Nottingham, 22nd– 29th August, 1866. Journal of the Statistical Society of London, 29(4):610–611, December 1866. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338652.

Anonymous:1866:PSS

[240] Anonymous. Proceedings of the Statistical Society. Journal of the Statistical Society of London, 29(4):612–613, December 1866. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338653.

Anonymous:1866:Md

[241] Anonymous. Miscellanea. Journal of the Statistical Society of London, 29(4): 614–632, December 1866. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338654.

Anonymous:1866:QRd

[242] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 29 (4):633–664, December 1866. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338655.

Anonymous:1867:VIa

[243] Anonymous. Volume information. Journal of the Statistical Society of London, 30(1):i–vii, March 1867. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338490.

Waley:1867:SCR

[244] Jacob Waley. On strikes and combinations, with reference to wages and the conditions of labour. Journal of the Statistical Society of London, 30(1):1– 20, March 1867. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338491.

Mouat:1867:PDS

[245] Frederic J. Mouat. On prison discipline and statistics in lower Bengal. Journal of the Statistical Society of London, 30(1): 21–57, March 1867. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338492.

Sykes:1867:FTB

[246] W. H. Sykes. Free trade in banking. Journal of the Statistical Society of London, 30(1):58–67, March 1867. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338493.

Worthington:1867:UPB

[247] A. W. Worthington. On the unequal proportion between the male and female population of some manufacturing and other towns. Journal of the Statistical Society of London, 30(1):68–79, March 1867. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338494.

Sargant:1867:PEE

[248] William Lucas Sargant. On the progress of elementary education. Journal of the Statistical Society of London, 30(1):80– 137, March 1867. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338495.

Anonymous:1867:Ma

[249] Anonymous. Miscellanea. Journal of the Statistical Society of London, 30(1):138– 177, March 1867. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338496.

Anonymous:1867:QRa

[250] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 30(1):178–210, March 1867. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338497.

Anonymous:1867:RCF

[251] Anonymous. Report of the Council for the financial year ended 31st December, 1866, and for the sessional year ended 15th March, 1867, presented at the thirty-third anniversary meeting of the Statistical Society, held at the Society’s rooms, 12, St. James’s Square, on Friday, 15th March, 1867; with the proceedings of that meeting. Journal of the Statistical Society of London, 30(2):211– 215, June 1867. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338511.

Balfour:1867:MCF

[252] Major-General Balfour. On the military conscription of France. Journal of the Statistical Society of London, 30(2):216– 292, June 1867. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338512.

Guy:1867:MLH

[253] William A. Guy. On the mortality of London hospitals: and incidentally on the deaths in the prisons and public institutions of the Metropolis. Journal of the Statistical Society of London, 30(2): 293–322, June 1867. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338513.

Newmarch:1867:CSC

[254] W. Thomas Newmarch. Central Statistical Commission of Austria; analysis of report for 1866. Journal of the Statistical Society of London, 30(2):323– 331, June 1867. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338514.

Anonymous:1867:Mb

[255] Anonymous. Miscellanea. Journal of the Statistical Society of London, 30(2):332– 345, June 1867. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338515.

Anonymous:1867:QRb

[256] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 30(2):346–373, June 1867. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338516.

Hammick:1867:JSE

[257] James T. Hammick. On the judicial statistics of England and Wales, with special reference to the recent returns relating to crime. Journal of the Statistical Society of London, 30(3):375–426, September 1867. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338841.

Bovill:1867:SCP

[258] William John Bovill. On the statistics of civil procedure in English courts of law. Journal of the Statistical Society of London, 30(3):427–453, September 1867. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338842.

Anonymous:1867:Mc

[259] Anonymous. Miscellanea. Journal of the Statistical Society of London, 30(3): 454–461, September 1867. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338843.

Anonymous:1867:QRc

[260] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 30 (3):462–493, September 1867. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338844.

Anonymous:1867:VIb

[261] Anonymous. Volume information. Journal of the Statistical Society of London, 30(4):605–620, December 1867. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338810.

Grant-Duff:1867:OAP

[262] M. E. Grant-Duff. Opening address of the President of Section F (Economic Science and Statistics), of the British Association for the Advancement of Science, at the thirty-seventh meeting, at Dundee, September, 1867. Journal of the Statistical Society of London, 30(4): 495–501, December 1867. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338811.

Clarke:1867:PIT

[263] Hyde Clarke. On public instruction in Turkey. Journal of the Statistical Society of London, 30(4):502–534, December 1867. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338812.

Sykes:1867:SSG

[264] W. H. Sykes. Statistics of sums given by native gentry in India for charitable and educational institutions and for works of public utility. Journal of the Statistical Society of London, 30(4):535–547, December 1867. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338813.

Renals:1867:AHT

[265] E. Renals. On arbitration in the hosiery trades of the Midland Counties. Journal of the Statistical Society of London, 30 (4):548–556, December 1867. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338814.

Rogers:1867:FAD

[266] J. E. Thorold Rogers. On the funds available for developing the machinery of education in England. Journal of the Statistical Society of London, 30(4): 557–561, December 1867. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338815.

Anonymous:1867:TSM

[267] Anonymous. Thirty-seventh meeting of the British Association for the Advancement of Science, held at Dundee, 4th– 11th September, 1867. Journal of the Statistical Society of London, 30(4):562– 563, December 1867. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338816.

Anonymous:1867:PSS

[268] Anonymous. Proceedings of the Statistical Society. Journal of the Statistical Society of London, 30(4):564–565, December 1867. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338817.

Anonymous:1867:Md

[269] Anonymous. Miscellanea. Journal of the Statistical Society of London, 30(4): 566–577, December 1867. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338818.

Anonymous:1867:QRd

[270] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 30 (4):578–604, December 1867. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338819.

Anonymous:1868:VIa

[271] Anonymous. Volume information. Journal of the Statistical Society of London, 31(1):i–vii, March 1868. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338880.

Heywood:1868:FGE

[272] James Heywood. Form of government and educational system of the University of Cambridge. Journal of the Statistical Society of London, 31(1):1–10, March 1868. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338881.

Brown:1868:RSI

[273] Samuel Brown. Report on the sixth International Statistical Congress, held at Florence, from 29th September to 5th October, 1867. Journal of the Statistical Society of London, 31(1):11–24, March 1868. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338882.

Hamilton:1868:TCR

[274] Archibald Hamilton. On the trade with the coloured races of Africa. Journal of the Statistical Society of London, 31(1): 25–48, March 1868. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338883.

Sykes:1868:ARU

[275] W. H. Sykes. Analysis of the report upon the state of the empire of France, presented to the Senate and legislative body, February, 1867. Journal of the Statistical Society of London, 31(1):49– 75, March 1868. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338884.

Anonymous:1868:Ma

[276] Anonymous. Miscellanea. Journal of the Statistical Society of London, 31(1):76– 99, March 1868. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338885.

Anonymous:1868:QRa

[277] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 31(1):100–120, March 1868. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338886.

Anonymous:1868:RCF

[278] Anonymous. Report of the Council for the financial year ended 31st December, 1867, and for the sessional year ended 15th March, 1868, presented at the thirty-fourth anniversary meeting of the Statistical Society, held at the Society’s rooms, 12, St. James’s Square, on Monday, 16th March, 1868; with the proceedings of that meeting. Journal of the Statistical Society of London, 31(2): 121–126, June 1868. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338857.

Caird:1868:ASU

[279] James Caird. On the agricultural statistics of the United Kingdom. Journal of the Statistical Society of London, 31(2): 127–145, June 1868. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338858.

Brown:1868:CPE

[280] Samuel Brown. On the comparative population of European states. Journal of the Statistical Society of London, 31(2):146–163, June 1868. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338859.

Courtney:1868:FUS

[281] Leonard H. Courtney. On the finances of the United States of America, 1861– 67. Journal of the Statistical Society of London, 31(2):164–221, June 1868. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338860.

Anonymous:1868:TCA

[282] Anonymous. Topographical classification of the agricultural statistics of England for 1867, illustrated. Journal of the Statistical Society of London, 31(2):222– 227, June 1868. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338861.

Anonymous:1868:Mb

[283] Anonymous. Miscellanea. Journal of the Statistical Society of London, 31(2):228– 238, June 1868. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338862.

Anonymous:1868:QRb

[284] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 31(2):239–263, June 1868. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338863.

Bikelas:1868:SKG

[285] Demetrius Bikelas. Statistics of the Kingdom of Greece. Journal of the Statistical Society of London, 31(3):265– 298, September 1868. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338853.

Elliott:1868:IMP

[286] J. H. Elliott. The increases of material prosperity and of moral agents, compared with the state of crime and pauperism. Journal of the Statistical Society of London, 31(3):299–330, September 1868. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338854.

Anonymous:1868:Mc

[287] Anonymous. Miscellanea. Journal of the Statistical Society of London, 31(3): 331–352, September 1868. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338855.

Anonymous:1868:QRc

[288] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 31 (3):353–372, September 1868. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338856.

Anonymous:1868:VIb

[289] Anonymous. Volume information. Journal of the Statistical Society of London, 31(4):503–519, December 1868. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338791.

Brown:1868:OAP

[290] Samuel Brown. Opening address of the President of Section F (Economic Science and Statistics), of the British Association for the Advancement of Science, at the thirty-eighth meeting, at Norwich, August, 1868. Journal of the Statistical Society of London, 31(4):373– 392, December 1868. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338792.

Heywood:1868:SHE

[291] James Heywood. State of health, and of the employment of Indians in the New England Company’s settlement of Kauyageh, Canada. Journal of the Statistical Society of London, 31(4):393– 394, December 1868. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338793.

Smith:1868:CPN

[292] William Smith. The cattle plague in Norfolk. Journal of the Statistical Society of London, 31(4):395–406, December 1868. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338794.

Mann:1868:SSR

[293] Horace Mann. Some statistics relating to the civil service. Journal of the Statistical Society of London, 31(4):407–417, December 1868. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338795.

Jeula:1868:PAS

[294] Henry Jeula. The present aspect of statistical inquiry in relation to shipping casualties. Journal of the Statistical Society of London, 31(4):418–425, December 1868. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338796.

Jevons:1868:CMC

[295] W. Stanley Jevons. On the condition of the metallic currency of the United Kingdom, with reference to the question of international coinage. Journal of the Statistical Society of London, 31 (4):426–464, December 1868. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338797.

Anonymous:1868:TEM

[296] Anonymous. Thirty-eighth meeting of the British Association for the Advancement of Science, held at Norwich, 19th– 26th August, 1868. Journal of the Statistical Society of London, 31(4):465–466, December 1868. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338798.

Anonymous:1868:PSS

[297] Anonymous. Proceedings of the Statistical Society. Journal of the Statistical Society of London, 31(4):467–468, December 1868. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338799.

Anonymous:1868:Md

[298] Anonymous. Miscellanea. Journal of the Statistical Society of London, 31(4):

469–482, December 1868. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338800.

Anonymous:1868:QRd

[299] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 31 (4):483–502, December 1868. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338801.

Anonymous:1869:VIa

[300] Anonymous. Volume information. Journal of the Statistical Society of London, 32(1):i–viii, March 1869. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338873.

Mann:1869:SNR

[301] Dr. Mann. Statistical notes regarding the Colony of Natal. Journal of the Statistical Society of London, 32(1):1– 28, March 1869. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338874.

Fielder:1869:RPF

[302] Charles Henry Fielder. On the rise, progress, and future prospects of tea cultivation in British India. Journal of the Statistical Society of London, 32(1):29– 37, March 1869. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338875.

Mann:1869:COC

[303] Horace Mann. On the cost and organization of the civil service. Journal of the Statistical Society of London, 32(1):38– 60, March 1869. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338876.

Caird:1869:ASU

[304] James Caird. On the agricultural statistics of the United Kingdom (second paper). Journal of the Statistical Society of London, 32(1):61–77, March 1869. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338877.

Anonymous:1869:Ma

[305] Anonymous. Miscellanea. Journal of the Statistical Society of London, 32(1):78– 124, March 1869. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338878.

Anonymous:1869:QRa

[306] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 32(1):125–151, March 1869. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338879.

Anonymous:1869:RCF

[307] Anonymous. Report of the Council for the financial year ended 31st December, 1868, and for the sessional year ended 15th March, 1869, presented at the thirty-fifth anniversary meeting of the Statistical Society, held at the Society’s rooms, 12, St. James’s Square, on Monday, 15th March, 1869; with the proceedings of that meeting. Journal of the Statistical Society of London, 32(2): 153–158, June 1869. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338820.

Guy:1869:ICP

[308] William A. Guy. On insanity and crime; and on the plea of insanity in criminal cases. Journal of the Statistical Society of London, 32(2):159–191, June 1869. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338821.

Brown:1869:SKN

[309] Samuel Brown. On the statistics of the Kingdom of The Netherlands. Journal of the Statistical Society of London, 32(2): 192–214, June 1869. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338822.

Anonymous:1869:Mb

[310] Anonymous. Miscellanea. Journal of the Statistical Society of London, 32(2):215– 233, June 1869. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338823.

Anonymous:1869:QRb

[311] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 32(2):234–257, June 1869. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338824.

Northcote:1869:OAP

[312] Stafford Northcote. Opening address of the President of Section F (Economic Science and Statistics), of the British Association for the Advancement of Science, at the thirty-ninth meeting, at Exeter, August, 1869. Journal of the Statistical Society of London, 32(3):259–270, September 1869. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338738.

Welton:1869:CPO

[313] T. A. Welton. On the classification of the people by occupations; and on other subjects connected with population statistics of England. Journal of the Statistical Society of London, 32(3):271–287, September 1869. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338739.

Glover:1869:DST

[314] John Glover. On the decline of shipbuilding on the Thames. Journal of the Statistical Society of London, 32(3): 288–292, September 1869. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338740.

Hamilton:1869:EPN

[315] Archibald Hamilton. On the economic progress of New Zealand. Journal of the Statistical Society of London, 32(3): 293–307, September 1869. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338741.

Purdy:1869:PTR

[316] Frederick Purdy. The pressure of taxation on real property. Journal of the Statistical Society of London, 32(3):308– 324, September 1869. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338742.

Anonymous:1869:Mc

[317] Anonymous. Miscellanea. Journal of the Statistical Society of London, 32(3): 325–338, September 1869. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338743.

Anonymous:1869:QRc

[318] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 32 (3):339–358, September 1869. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338744.

Anonymous:1869:VIb

[319] Anonymous. Volume information. Journal of the Statistical Society of London, 32(4):485–502, December 1869. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338750.

Newmarch:1869:IAP

[320] William Newmarch. Inaugural address on the progress and present condition of statistical inquiry, delivered at the Society’s rooms, 12, St. James’s Square, London, on Tuesday, 16th November, 1869. Journal of the Statistical Society of London, 32(4):359–390, December 1869. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338751.

Brown:1869:RSI

[321] Samuel Brown. Report on the seventh International Statistical Congress, held at The Hague, 6th–11th September, 1869. Journal of the Statistical Society of London, 32(4):391–410, December 1869. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338752.

Palgrave:1869:HAE

[322] Robert Harry Inglis Palgrave. On the house accommodation of England and Wales, with reference to the census of 1871. Journal of the Statistical Society of London, 32(4):411–427, December 1869. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338753.

Helm:1869:CTU

[323] Elijah Helm. The cotton trade of the United Kingdom, during the seven years, 1862–68, as compared with the seven years, 1855–61; with remarks on the return of factories existing in 1868. Journal of the Statistical Society of London, 32(4):428–437, December 1869. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338754.

Anonymous:1869:Md

[324] Anonymous. Miscellanea. Journal of the Statistical Society of London, 32(4): 438–463, December 1869. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338755.

Anonymous:1869:QRd

[325] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 32 (4):464–483, December 1869. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338756.